100W and 500W bulbs are connected in series to mains which bulb will glow brighter

Grade:

1 Answers

Divyansh Nema
5 years ago
   
    POWER =  (CURRENT)^2  ×  RESIST
   
    P = I^2 R 
   
   IN SERIES CURRENT IS CONSTANT 
 
   100 / constant  =  R1 
   500 / constant  =  R2 
 
    R2 > R1 
    
    
 
    LET  V =  X  FOR BOTH BULBS 
    
   P  =  V^2 / R
    
   P1 = X^2 / R1           P2 = X^2 / R2
  
   P1 > P2
  
  THUS  100 W BULB GLOWS BRIGTER